OpenStudy (anonymous):

An electron has a speed of 300 m/s accurate to 0.01%. With what fundamental accuracy can we locate the position of the electron

OpenStudy (anonymous):

Use Heisenberg Uncertainty principle.

OpenStudy (anonymous):

@@ardalan deltax=h/mdeltaV ..... how r we gonna substitute ....

OpenStudy (anonymous):

use the mass of electron.

OpenStudy (anonymous):

can u pl give the formula.... mass is not given ..rt ....

OpenStudy (anonymous):

the mass of electron is :9.10938291 × 10-31 kilograms Planck's constant = 6.62606957 × 10-34 m2 kg / s you have delta V. then you will find delta x.

OpenStudy (anonymous):

kk well h is the planks constant ... them m remains ....deltaV=300 h= 6.62606957 × 10^-34 m2 kg / s whats m =?

OpenStudy (anonymous):

m is the mass of the particle. here is the mass of electron that i gave you in my last post. :)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

what does it mean? "kk" and "rt"

OpenStudy (chmvijay):

okok and right

OpenStudy (anonymous):

Oh! thanks! :)

OpenStudy (anonymous):

thanku ... got it...^_^